A 2.5m tall steel has a cross sectional area of 1.5m2. At the bottom with a height of 0.5m is liquid  water on top of which is a 1m high layer of gasoline. The gasoline surface is exposed to atmospheric  air at 101 kPa. What is the highest pressure in the water?


1
Expert's answer
2021-10-10T15:56:14-0400

"p=p_0+\rho_1 gh_1+\rho_2 gh_2,"

"p=101000+9.8\cdot (570\cdot 1+1000\cdot 0.5)=113.3~\text{kPa}."